# Proto-CLIP: Vision-Language Prototypical Network for Few-Shot Learning

### v1

@@ Abstract
We propose a novel framework for few-shot learning by leveraging large-scale vision-language models such as CLIP [1]. Motivated by unimodal prototypical networks for few-shot learning, we introduce Proto-CLIP that utilizes image prototypes and text prototypes for few-shot learning. Specifically, Proto-CLIP adapts the image encoder and text encoder in CLIP in a joint fashion using few-shot examples. The two encoders are used to compute prototypes of image classes for classification. During adaptation, we propose aligning the image prototypes and the text prototypes of the corresponding classes. Such alignment is beneficial for few-shot classification due to the reinforced contributions from both types of prototypes. We demonstrate the effectiveness of our method by conducting experiments on benchmark datasets for few-shot learning, as well as in the real world for robot perception [^1] Project page is available at https://irvlutd.github.io/Proto-CLIP.
Keywords: Robot Perception, Object Recognition, Few-Shot Learning, Contrastive-Learning, Vision-Language, Multimodal
1 Introduction
Building autonomous robots that can help people perform various tasks is the dream of every roboticist. Nowadays, most robots are working in factories and warehouses by performing pre-programmed repetitive tasks such as assembling and delivering. In the future, we believe that there will be intelligent robots that can perform tasks in human environments autonomously. For example, people can instruct a robot by saying “bring me a bottle of water” or “wash the mug on the table”, then the robot will execute the instructions accordingly. In these scenarios, robots need to recognize objects from sensory data in order to understand the required tasks. In this work, we develop a novel few-shot learning method that can enable robots to recognize novel objects from just a few example images per object.
@@ 1 Introduction
We believe that few-shot learning [2] is a promising paradigm to enable robots to recognize a large number of objects. The appeal lies in the ease of data collection—just a few example images is sufficient for teaching a robot a novel object. On the contrary, object model-based approaches build 3D models of objects and then use these 3D models [3] for object recognition. Object category-based approaches focus on recognizing category labels of objects such as 80 categories in the MSCOCO dataset [4]. The limitation of model-based object recognition is the difficulty of obtaining a large number of 3D models for many objects in the real world. Current 3D scanning techniques cannot deal well with metal objects or transparent objects. For category-based object recognition, it is difficult to obtain a large number of images for each category in robotic settings. Large-scale datasets for object categories such as ImageNet [5] and Visual Genome [6] are collected from the Internet. These Internet images are not very suitable for learning object representations for robot manipulation due to domain differences. Due to the limitations of model-based and category-based object recognition, if a robot can learn to recognize a new object from a few images of the object, it is likely to scale up the number of objects that the robot can recognize in the real world.
@@
Figure 1: Our Proto-CLIP model learns a joint embedding space of images and text, where image prototypes and text prototypes are learned and aligned using support sets for few-shot classification.
@@
2 Related Work
Method Use Support Sets Adapt Image Embedding Adapt Text Embedding Align Image and Text Zero-shot CLIP [1] Linear-probe CLIP [1] CoOp [8] CLIP-Adapter [9] Tip-Adapter [10] Proto-CLIP (Ours)
Table 1: Comparison between our proposed method with existing CLIP-based methods for few-shot learning. “Use Support Sets” indicates if a method uses support training sets for fine-tuning. “Adapt Image/Text Embedding” indicates if a method adapts the image/text embeddings in CLIP. “Align Image and Text” indicates if a method aligns images and text in the feature space.
@@ 2 Related Work
CLIP-based Few-Shot Learning. The CLIP [1] model applies contrastive learning to image-text pairs from the Internet. It consists of an image encoder and a text encoder for the extraction of features from images and text, respectively. Its training objective is to maximize the similarity between the corresponding image and text in a pair in a high-dimensional joint feature space. After training, CLIP can be used for zero-shot image classification by comparing image features with text embeddings of novel class names. This model is denoted as zero-shot CLIP. When a few training images are available for each class, several approaches are proposed to improve zero-shot CLIP. The linear-probe CLIP model [1] trains a logistic regression classifier using CLIP image features. CoOp [8] proposes to use learnable vectors as a prompt for the CLIP text encoder for few-shot learning. CLIP-Adapter [9] learns two layers of linear transformations on top of the image encoder and the text encoder with residual connections, respectively, to adapt CLIP features for few-shot learning. Tip-Adapter [10] builds a key-value cache model, where keys are CLIP image features and values are one-hot vectors of the class labels. Given a query image, its image feature is compared with the cache keys to combine the value labels for classification. Tip-Adapter can also fine-tune the keys by treating them as learnable parameters, which further improves the few-shot classification accuracy.
Table 1 compares our proposed method with existing CLIP-model-based few-shot learning methods. By using the image prototypes and text prototypes for classification, our method can adapt both the image embeddings and text embeddings from CLIP. In addition, the model aligns the image prototypes and the text prototypes, which serves as a regularization term in adapting the feature embeddings. We empirically verify our model by conducting experiments on benchmark datasets for few-shot learning.
@@ 3 Method
Figure 2: Overview of our proposed Proto-CLIP model. The CLIP image encoder and text encoder are frozen during training. The image memory, the text memory and the adapter network are learned. Given a class name, $\tau_{i}$ returns the $i^{th}$ out of $\tilde{K}$ predefined text prompts.
@@
$\mathbf{c}_{k}^{x}=\frac{1}{M_{k}}\sum_{y_{i}^{s}=k}\phi_{\text{Image}}(\mathbf{x}_{i}^{s}),\;\;\mathbf{c}_{k}^{y}=\frac{1}{\tilde{M_{k}}}\sum_{j=1}^{\tilde{M_{k}}}\phi_{\text{Text}}(\text{Prompt}_{j}(y_{i}^{s}=k)),$ (4)
@@
where $M_{k}$ is the number of examples with label $k$, and $\tilde{M_{k}}$ is the number of prompts for label $k$. To compute text embeddings, we can either directly input the class names such as “mug” and “plate” into the text encoder, or convert the class names to phases such as “a photo of mug” and “a photo of plate” and then input the phrases into the text encoder. These phrases are known as prompts of the vision-language models. We can use multiple prompts for each class label. $\phi_{\text{Image}}(\mathbf{x}_{i}^{s})$ and $\phi_{\text{Text}}(\text{Prompt}_{j}(y_{i}^{s}=k))$ denote the image embedding and the text embedding of the image-label pair $(\mathbf{x}_{i}^{s},y_{i}^{s})$ computed using the CLIP image encoder and the text encoder, respectively. These embeddings with dimension $C$ of the support set form the image memory and the text memory, as shown in Fig. 2. They are learnable embedding vectors initialized by the computed embeddings using the CLIP image encoder and text encoder. We use $\mathbf{c}_{k}^{x}$ and $\mathbf{c}_{k}^{y}$ to denote the mean of the embeddings of the images and the prompts for class $k$, respectively. Since the image embeddings and the text embeddings are of the same dimension, we can compute the distance between the text prototype $\mathbf{c}_{k}^{y}$ and the image embedding $g_{\mathbf{w}_{1}}(\mathbf{x}^{q})$ in Eq. (3). As we can see, our model leverages prototypical networks with image encoder and text encoder from CLIP. We name it “Proto-CLIP”.
@@
Figure 3: Two designs of the adapters. (a) A Multi-layer perceptron-based adapter as in [9]. (b) A convolution-based adapter we introduce. The feature dimension is for CLIP ResNet50 backbone.
@@
Learning the memories and the adapter. During training, we can construct a support set $\mathcal{S}=\{\mathbf{x}_{i}^{s},y_{i}^{s}\}_{i=1}^{M}$ and a query set with ground truth labels $\mathcal{Q}=\{\mathbf{x}_{j}^{q},y_{j}^{q}\}_{j=1}^{L}$. Then we can use $\mathcal{S}$ and $\mathcal{Q}$ to learn the weights in Proto-CLIP. First, the support set is used to initialize the image memory $\mathbf{W}_{\text{image}}$ and the text memory $\mathbf{W}_{\text{text}}$. Second, the weights in the adapter network applied to the query images $g_{\mathbf{w}_{1}}(\cdot)$ need to be learned. Fig. 3 shows two designs of the adapter network, i.e., an MLP-based adapter as in [9] and a convolution-based adapter we introduce. The convolution-based adapter has fewer weights to learn compared to the MLP-based one. We found that the two adapters have their own advantages on different datasets in our experiments. Finally, motivated by the CLIP-Adapter [9], we do not fine-tune the weights in the image encoder and text encoder by freezing these weights during training. In this way, we can reuse the weights of CLIP trained on a large number of image-text pairs and adapt the image embeddings and the text embeddings.
@@
Loss Functions. The first loss function is the negative log-probability of the true label for a query image: $\mathcal{L}_{1}(\mathbf{W}_{\text{image}},\mathbf{W}_{\text{text}},\mathbf{w}_{1})=-\log P(y^{q}=k|\mathbf{x}^{q},\mathcal{S})$, where $P(y^{q}=k|\mathbf{x}^{q},\mathcal{S})$ is defined in Eq. (1). Minimizing $\mathcal{L}_{1}$ learns the weights to classify the query images correctly. Second, we propose aligning the image prototypes and the text prototypes in training. Let $\{\mathbf{c}_{1}^{x},\mathbf{c}_{2}^{x},\ldots,\mathbf{c}_{N}^{x}\}$ be the image prototypes computed from the image embeddings for the $N$ classes and $\{\mathbf{c}_{1}^{y},\mathbf{c}_{2}^{y},\ldots,\mathbf{c}_{N}^{y}\}$ be the corresponding text prototypes. We would like to learn the model weights such that $\mathbf{c}_{k}^{x}$ is close to $\mathbf{c}_{k}^{y}$ and far from other prototypes in the embedding space. We utilize the InfoNCE loss for contrastive learning [21]:
$\mathcal{L}_{2}^{k}(\mathbf{c}_{k}^{x},\{\mathbf{c}_{k^{\prime}}^{y}\}_{k^{\prime}=1}^{N})=-\log\frac{\exp(\mathbf{c}_{k}^{x}\cdot\mathbf{c}_{k}^{y})}{\sum_{k^{\prime}=1}^{N}\exp(\mathbf{c}_{k}^{x}\cdot\mathbf{c}_{k^{\prime}}^{y})},\mathcal{L}_{3}^{k}(\mathbf{c}_{k}^{y},\{\mathbf{c}_{k^{\prime}}^{x}\}_{k^{\prime}=1}^{N})=-\log\frac{\exp(\mathbf{c}_{k}^{y}\cdot\mathbf{c}_{k}^{x})}{\sum_{k^{\prime}=1}^{N}\exp(\mathbf{c}_{k}^{y}\cdot\mathbf{c}_{k^{\prime}}^{x})}$ (5)
@@
for $k=1,\dots,N$, where $\cdot$ indicates dot-product. Here, $\mathcal{L}_{2}^{k}(\mathbf{c}_{k}^{x},\{\mathbf{c}_{k^{\prime}}^{y}\}_{k^{\prime}=1}^{N})$ compares an image prototype $\mathbf{c}_{k}^{x}$ with the text prototypes $\{\mathbf{c}_{k^{\prime}}^{y}\}_{k^{\prime}=1}^{N}$, while $\mathcal{L}_{3}^{k}(\mathbf{c}_{k}^{y},\{\mathbf{c}_{k^{\prime}}^{x}\}_{k^{\prime}=1}^{N})$ compares a text prototype $\mathbf{c}_{k}^{y}$ with the image prototypes $\{\mathbf{c}_{k^{\prime}}^{x}\}_{k^{\prime}=1}^{N}$. In this way, we can align the image prototypes and the text prototypes for the $N$ classes. This alignment can facilitate classification, since the class conditional probabilities are computed using the image prototypes and the text prototypes as in Eqs. (2) and (3). The total loss function for training is:
$\mathcal{L}=-\frac{1}{L}\sum_{j=1}^{L}\log P(y_{j}^{q}=k|\mathbf{x}_{j}^{q},\mathcal{S})+\frac{1}{N}\sum_{k=1}^{N}\big{(}\mathcal{L}_{2}^{k}(\mathbf{c}_{k}^{x},\{\mathbf{c}_{k^{\prime}}^{y}\}_{k^{\prime}=1}^{N})+\mathcal{L}_{3}^{k}(\mathbf{c}_{k}^{y},\{\mathbf{c}_{k^{\prime}}^{x}\}_{k^{\prime}=1}^{N})\big{)}$ (6)
@@
for a query set $\mathcal{Q}=\{\mathbf{x}_{j}^{q},y_{j}^{q}\}_{j=1}^{L}$. Following previous CLIP-based few-shot learning methods [8, 9, 10], the support set and the query set are the same during training in our experiments, i.e., $\mathcal{S}=\mathcal{Q}$.
4 Experiments
Datasets and Evaluation Metric. Following previous CLIP-based few-shot learning methods [8, 9, 10], we conduct experiments on the following datasets for evaluation: ImageNet [5], StandfordCars [22], UCF101 [23], Caltech101 [24], Flowers102 [25], SUN397 [26], DTD [27], EuroSAT [28], FGVCAircraft [29], OxfordPets [30], and Food101 [31]. In addition, we also include the FewSOL dataset [15] recently introduced for few-shot object recognition in robotic environments. In the $N$-way $K$-shot classification setting, $K$ images for each class will be sampled from each dataset for training. A validation set of each dataset is reserved for hyper-parameter tuning, and a test set is used for evaluation. We report the classification accuracy of the test set as an evaluation metric.
Choosing the Hyper-parameters: $\alpha$ and $\beta$. From the experiments, we found that the two hyper-parameters $\alpha$ in Eq. (1) and $\beta$ in Eqs. (2) and (3) play a critical role in classification accuracy. Therefore, for each dataset, we conducted a grid search of the two parameters using the validation set. Then we finalize their values for all the runs in our experiments.
Proto-CLIP Variants. i) “Proto-CLIP”: we do not train the image memory and the text memory and do not use any adapter in Proto-CLIP (Fig. 2), we directly run inference using the pre-trained CLIP features. ii) “Proto-CLIP-$F$”: we train the image memory and/or the text memory with the adapter. During training, for all the query images, we precompute their CLIP image features and directly use these stored features for training. This variant can be trained more quickly. Therefore, we use it for our ablation studies. iii) “Proto-CLIP-$F$-$Q^{T}$”: During training, for each query image, we apply random data augmentation operations such as cropping and horizontal flip. Then we need to compute CLIP image features for the transformed query images during training.
@@ 4 Experiments
Figure 4: Barnes-Hut t-SNE visualization [32] using the FewSOL dataset [15]. (a) Image and text prototypes from zero-shot CLIP, which are not aligned. (b) Aligned image and text prototypes from fine-tuned Proto-CLIP.
4.1 Ablation Studies

### v3

@@ Abstract
We propose a novel framework for few-shot learning by leveraging large-scale vision-language models such as CLIP [1]. Motivated by unimodal prototypical networks for few-shot learning, we introduce Proto-CLIP which utilizes image prototypes and text prototypes for few-shot learning. Specifically, Proto-CLIP adapts the image and text encoder embeddings from CLIP in a joint fashion using few-shot examples. The embeddings from the two encoders are used to compute the respective prototypes of image classes for classification. During adaptation, we propose aligning the image and text prototypes of the corresponding classes. Such alignment is beneficial for few-shot classification due to the reinforced contributions from both types of prototypes. Proto-CLIP has both training-free and fine-tuned variants. We demonstrate the effectiveness of our method by conducting experiments on benchmark datasets for few-shot learning, as well as in the real world for robot perception [^1] Project page: https://irvlutd.github.io/Proto-CLIP.
I INTRODUCTION
@@
Building autonomous robots that can help people perform various tasks is the dream of every roboticist. Nowadays, most robots are working in factories and warehouses by performing pre-programmed repetitive tasks such as assembling and delivering. In the future, we believe that there will be intelligent robots that can perform tasks in human environments autonomously. For example, people can instruct a robot by saying “bring me a bottle of water” or “wash the mug on the table”, then the robot will execute the instructions accordingly. In these scenarios, robots need to recognize objects from sensory data in order to understand the required tasks. In this work, we develop a novel few-shot learning method that can enable robots to recognize novel objects from just a few example images per object. We believe that few-shot learning [2] is a promising paradigm to enable robots to recognize a large number of objects. The appeal lies in the ease of data collection—just a few example images is sufficient for teaching a robot a novel object. On the contrary, object model-based approaches build 3D models of objects and then use these 3D models [3] for object recognition. Object category-based approaches focus on recognizing category labels of objects such as 80 categories in the MSCOCO dataset [4]. The limitation of model-based object recognition is the difficulty of obtaining a large number of 3D models for many objects in the real world. Current 3D scanning techniques cannot deal well with metal objects or transparent objects. For category-based object recognition, it is difficult to obtain a large number of images for each category in robotic settings. Large-scale datasets for object categories such as ImageNet [5] and Visual Genome [6] are collected from the Internet. These Internet images are not very suitable for learning object representations for robot manipulation due to domain differences. Due to the limitations of model-based and category-based object recognition, if a robot can learn to recognize a new object from a few images of the object, it is likely to scale up the number of objects that the robot can recognize in the real world.
@@
Figure 1: Our Proto-CLIP model learns a joint embedding space of images and text, where image and text prototypes formed using support sets are learned and aligned for few-shot classification.
@@
II RELATED WORK
Method Use Support Sets Adapt Image Embedding Adapt Text Embedding Align Image and Text
Zero-shot CLIP [1]
Linear-probe CLIP [1]
CoOp [8]
CLIP-Adapter [9]
Tip-Adapter [10]
Sus-X [16]
Proto-CLIP (Ours)
TABLE I: Comparison of our proposed method with the existing CLIP-based few-shot learning methods. “Use Support Sets” indicates if a method uses support training sets for fine-tuning. “Adapt Image/Text Embedding” indicates if a method adapts the image/text embeddings obtained from CLIP. “Align Image and Text” indicates if a method specifically aligns images and text in the feature space.
@@ II RELATED WORK
CLIP-based Few-Shot Learning. The CLIP [1] model applies contrastive learning to image-text pairs from the Internet. It consists of an image encoder and a text encoder for the extraction of features from images and text, respectively. Its training objective is to maximize the similarity between the corresponding image and text in a pair in a high-dimensional joint feature space. After training, CLIP can be used for zero-shot image classification by comparing image features with text embeddings of novel class names. This model is denoted as zero-shot CLIP. When a few training images are available for each class, several approaches are proposed to improve zero-shot CLIP. The linear-probe CLIP model [1] trains a logistic regression classifier using CLIP image features. CoOp [8] proposes to use learnable vectors as a prompt for the CLIP text encoder for few-shot learning. CLIP-Adapter [9] learns two layers of linear transformations on top of the image encoder and the text encoder with residual connections, respectively, to adapt CLIP features for few-shot learning. Tip-Adapter [10] builds a key-value cache model, where keys are CLIP image features and values are one-hot vectors of the class labels. Given a query image, its image feature is compared with the cache keys to combine the value labels for classification. Tip-Adapter can also fine-tune the keys by treating them as learnable parameters, which further improves the few-shot classification accuracy. Sus-X [16] leverages the power of Stable Diffusion [17] to create support sets and aims to address the issue of uncalibrated intra-modal embedding distances in TIP-Adapter [10] by utilizing inter-modal distances as a connecting mechanism. Table I compares our proposed method with existing CLIP-model-based few-shot learning methods. By using the image prototypes and text prototypes for classification, our method can adapt both the image embeddings and text embeddings from CLIP. In addition, the model aligns the image prototypes and the text prototypes, which serves as a regularization term in adapting the feature embeddings. We empirically verify our model by conducting experiments on benchmark datasets for few-shot learning.
@@ III METHOD
Figure 2: Overview of our proposed Proto-CLIP model. The image memory, the text memory and the adapter network are learned. Given a class name, $\tau_{i}$ returns the $i^{th}$ out of $\tilde{K}$ predefined text prompts.
@@
$\mathbf{c}_{k}^{x}=\frac{1}{M_{k}}\sum_{y_{i}^{s}=k}\phi_{\text{Image}}(\mathbf{x}_{i}^{s})$ (4)
$\;\;\mathbf{c}_{k}^{y}=\frac{1}{\tilde{M_{k}}}\sum_{j=1}^{\tilde{M_{k}}}\phi_{\text{Text}}(\text{Prompt}_{j}(y_{i}^{s}=k)),$ (5)
@@
where $M_{k}$ is the number of examples with label $k$, and $\tilde{M_{k}}$ is the number of prompts for label $k$. To compute text embeddings, we can either directly input the class names such as “mug” and “plate” into the text encoder, or convert the class names to phrases such as “a photo of mug” and “a photo of plate” and then input the phrases into the text encoder. These phrases are known as prompts of the vision-language models. We can use multiple prompts for each class label. $\phi_{\text{Image}}(\mathbf{x}_{i}^{s})$ and $\phi_{\text{Text}}(\text{Prompt}_{j}(y_{i}^{s}=k))$ denote the image embedding and the $j$th text embedding of the image-label pair $(\mathbf{x}_{i}^{s},y_{i}^{s})$ computed using the CLIP image encoder and the text encoder, respectively. These embeddings with dimension $C$ of the support set form the image memory and the text memory, as shown in Fig. 2. They are learnable embedding vectors initialized by the computed embeddings using the CLIP image encoder and text encoder. We use $\mathbf{c}_{k}^{x}$ and $\mathbf{c}_{k}^{y}$ to denote the mean of the embeddings of the images and the prompts for class $k$, respectively. Since the image embeddings and the text embeddings are of the same dimension, we can compute the distance between the text prototype $\mathbf{c}_{k}^{y}$ and the image embedding $g_{\mathbf{w}_{1}}(\mathbf{x}^{q})$ in Eq. 3. As we can see, our model leverages prototypical networks with image encoder and text encoder from CLIP. We name it “Proto-CLIP”.
@@
Figure 3: Two designs of the adapters. (a) A Multi-layer perceptron-based adapter as in [9]. (b) A convolution-based adapter that we introduce. The feature dimension is for CLIP ResNet50 backbone.
@@
Learning the memories and the adapter. During training, we can construct a support set $\mathcal{S}=\{\mathbf{x}_{i}^{s},y_{i}^{s}\}_{i=1}^{M}$ and a query set with ground truth labels $\mathcal{Q}=\{\mathbf{x}_{j}^{q},y_{j}^{q}\}_{j=1}^{L}$. Then we can use $\mathcal{S}$ and $\mathcal{Q}$ to learn the weights in Proto-CLIP. First, the support set is used to initialize the image memory $\mathbf{W}_{\text{image}}$ and the text memory $\mathbf{W}_{\text{text}}$. Second, the weights in the adapter network applied to the query images $g_{\mathbf{w}_{1}}(\cdot)$ need to be learned. Fig. 3 shows two designs of the adapter network, i.e., an MLP-based adapter as in [9] and a convolution-based adapter that we introduce. The convolution-based adapter has fewer weights to learn compared to the MLP-based one. We found that the two adapters have their own advantages on different datasets in our experiments. Finally, motivated by the CLIP-Adapter [9], we do not fine-tune the weights in the image encoder and text encoder by freezing these weights during training. In this way, we can reuse the weights of CLIP trained on a large number of image-text pairs and adapt the image embeddings and the text embeddings.
@@
Loss Functions. The first loss function is the negative log-probability of the true label for a query image: $\mathcal{L}_{1}(\mathbf{W}_{\text{image}},\mathbf{W}_{\text{text}},\mathbf{w}_{1})=-\log P(y^{q}=k|\mathbf{x}^{q},\mathcal{S})$, where $P(y^{q}=k|\mathbf{x}^{q},\mathcal{S})$ is defined in Eq. 1. Minimizing $\mathcal{L}_{1}$ learns the weights to classify the query images correctly. Second, we propose aligning the image prototypes and the text prototypes in training. Let $\{\mathbf{c}_{1}^{x},\mathbf{c}_{2}^{x},\ldots,\mathbf{c}_{N}^{x}\}$ be the image prototypes computed from the image embeddings for the $N$ classes and $\{\mathbf{c}_{1}^{y},\mathbf{c}_{2}^{y},\ldots,\mathbf{c}_{N}^{y}\}$ be the corresponding text prototypes. We would like to learn the model weights such that $\mathbf{c}_{k}^{x}$ is close to $\mathbf{c}_{k}^{y}$ and far from other prototypes in the embedding space. We utilize the InfoNCE loss for contrastive learning [23]:
$\mathcal{L}_{2}^{k}(\mathbf{c}_{k}^{x},\{\mathbf{c}_{k^{\prime}}^{y}\}_{k^{\prime}=1}^{N})=-\log\frac{\exp(\mathbf{c}_{k}^{x}\cdot\mathbf{c}_{k}^{y})}{\sum_{k^{\prime}=1}^{N}\exp(\mathbf{c}_{k}^{x}\cdot\mathbf{c}_{k^{\prime}}^{y})}$ (6)
$\mathcal{L}_{3}^{k}(\mathbf{c}_{k}^{y},\{\mathbf{c}_{k^{\prime}}^{x}\}_{k^{\prime}=1}^{N})=-\log\frac{\exp(\mathbf{c}_{k}^{y}\cdot\mathbf{c}_{k}^{x})}{\sum_{k^{\prime}=1}^{N}\exp(\mathbf{c}_{k}^{y}\cdot\mathbf{c}_{k^{\prime}}^{x})}$ (7)
@@
for $k=1,\dots,N$, where $\cdot$ indicates dot-product. Here, $\mathcal{L}_{2}^{k}(\mathbf{c}_{k}^{x},\{\mathbf{c}_{k^{\prime}}^{y}\}_{k^{\prime}=1}^{N})$ compares an image prototype $\mathbf{c}_{k}^{x}$ with the text prototypes $\{\mathbf{c}_{k^{\prime}}^{y}\}_{k^{\prime}=1}^{N}$, while $\mathcal{L}_{3}^{k}(\mathbf{c}_{k}^{y},\{\mathbf{c}_{k^{\prime}}^{x}\}_{k^{\prime}=1}^{N})$ compares a text prototype $\mathbf{c}_{k}^{y}$ with the image prototypes $\{\mathbf{c}_{k^{\prime}}^{x}\}_{k^{\prime}=1}^{N}$. In this way, we can align the image prototypes and the text prototypes for the $N$ classes. This alignment can facilitate classification, since the class conditional probabilities are computed using the image prototypes and the text prototypes as in Eqs. 2 and 3. The total loss function for training is:
$\begin{split}\mathcal{L}&=-\frac{1}{L}\sum_{j=1}^{L}\log P(y_{j}^{q}=k|\mathbf{x}_{j}^{q},\mathcal{S})\\ &+\frac{1}{N}\sum_{k=1}^{N}\big{(}\mathcal{L}_{2}^{k}(\mathbf{c}_{k}^{x},\{\mathbf{c}_{k^{\prime}}^{y}\}_{k^{\prime}=1}^{N})+\mathcal{L}_{3}^{k}(\mathbf{c}_{k}^{y},\{\mathbf{c}_{k^{\prime}}^{x}\}_{k^{\prime}=1}^{N})\big{)}\end{split}$ (8)
@@
for a query set $\mathcal{Q}=\{\mathbf{x}_{j}^{q},y_{j}^{q}\}_{j=1}^{L}$. Following previous CLIP-based few-shot learning methods [8, 9, 10], the support set and the query set are the same during training in our experiments, i.e., $\mathcal{S}=\mathcal{Q}$ meaning any of the support samples can act as a query sample during training.
@@
Figure 4: Barnes-Hut t-SNE visualization [24] using the FewSOL dataset [15]. (a) Image and text prototypes from zero-shot CLIP, which are not aligned. (b) Aligned image and text prototypes from Proto-CLIP-$F$.
IV EXPERIMENTS
Datasets and Evaluation Metric. Following previous CLIP-based few-shot learning methods [8, 9, 10], we conduct experiments on the following datasets for evaluation: ImageNet [5], StandfordCars [25], UCF101 [26], Caltech101 [27], Flowers102 [28], SUN397 [29], DTD [30], EuroSAT [31], FGVCAircraft [32], OxfordPets [33], and Food101 [34]. In addition, we also include the FewSOL dataset [15] recently introduced for few-shot object recognition in robotic environments in order to improve object classification for robot manipulation tasks. In the $N$-way $K$-shot classification setting, $K$ images for each class will be sampled from each dataset for training. A validation set of each dataset is reserved for hyper-parameter tuning, and a test set is used for evaluation. We evaluate using test set classification accuracy, as in related works.
Choosing the Hyper-parameters: $\alpha$ and $\beta$. From the experiments, we found that the two hyper-parameters $\alpha$ in Eq. 1 and $\beta$ in Eq. 2 and Eq. 3 play a critical role in classification accuracy. Therefore, for each dataset, we conducted a grid search of the two parameters using the validation set. Then we finalize their values for all the runs in our experiments.
Proto-CLIP Variants. i) “Proto-CLIP”: we do not train the image memory and the text memory and do not use any adapter in Proto-CLIP (Fig. 2), we directly run inference using the pre-trained CLIP features. We term this variant the “training-free” version because it does not require training. This offers a convenient way to quickly test new datasets without the complexities of training, although it comes with the caveat of potential misalignment between visual and textual features. ii) “Proto-CLIP-$F$”: we train the image memory and/or the text memory with the adapter. During training, for all the query images, we precompute their CLIP image features and directly use these stored features for training. This variant can be trained more quickly w.r.t. the following variant. Therefore, we use it for our ablation studies. iii) “Proto-CLIP-$F$-$Q^{T}$”: During training, for each query image, we apply random data augmentation operations such as cropping and horizontal flip. Then we compute CLIP image features for the transformed query images during training.
IV-A Ablation Studies